OpenStudy (anonymous):

Confused in this problem :P When the digits of a two digit number are reversed, the new number is 36 more than the original number. The units digit is twice the tens digit. What is the original number?

OpenStudy (anonymous):

Let digits of original number be x and y. 10x + y + 36 = 10y + x It also says y = 2x. Substitute and solve.

OpenStudy (anonymous):

O so thats how you set it up...

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Is this correct y=4?

OpenStudy (anonymous):

10x + 2x + 36 = 20x + x => 12x + 36 = 21x => 9x = 36 => x = 4 => y = 8

OpenStudy (anonymous):

Can I ask you one question?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

For the part "36 more than the original number" what if it said 36 LESS THAN would it be set up like this 10x+2x-36?

OpenStudy (anonymous):

10x + y - 36 = 10y + x

OpenStudy (anonymous):

You just need to think about how to formulate. You know x is the tens digit and y is the ones digit in the original number. So, the original number will be 10x + y. The "Flipped" number will be 10y + x. Now, you need to include the "relation" between original number and flipped number. Flipped number is 36 more than original? => 10y + x = 10x + y + 36 Flipped number is half the original number? => 10y + x = (10x + y)/2 Flipped number is 10 less than the original number? => 10y + x = 10x + y - 10 Etc....!

OpenStudy (anonymous):

How do you set up if the sum of the digits is 8?

OpenStudy (anonymous):

Like this 10x+y=8?

OpenStudy (anonymous):

If sum of the digits is 8, then you simply have x+y = 8 because that is just the simple addition of the digits. Not the "value" of the digits.

OpenStudy (anonymous):

Ok then.

OpenStudy (anonymous):

Remember, the "number" assigned to a digit is NOT the same as the "value" of that digit.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

In the number 386, the tens digit is 8. But, its "value" is 80.

OpenStudy (anonymous):

Hundreds digit is 3 and its value is 300.

OpenStudy (anonymous):

Ok